OpenStudy (anonymous):

What is the result of isolating y^2 in the equation below? (x+6)^2 + y^2 = 1 A.) y^2 = -x^2 - 12x -35 B.) y^2 = x^2 + 12x + 37 C.) y^2 = -x^2 - 35 D.) y^2 = x^2 + 37 Help? Please?

OpenStudy (anonymous):

(x+6)^2 + y^2 = 1 First, simplify. x^2 + 12x + 36 + y^2 = 1 Next subtract x^2 from both side of the equation. 36 + 12x + y^2 = -x^2 +1 Next subtract 12x from both side of the equation. 36+ y^2 = -x^2 -12x Next subtract 36 from both side of the equation. y^2 = -x^2 -12x -35

OpenStudy (anonymous):

Okay thanks.

OpenStudy (anonymous):

No problem. :)